OpenStudy (timaashorty):

Urgent algebra help ;c Write the 10th term of the sequence.

OpenStudy (timaashorty):

okay

OpenStudy (jdoe0001):

what's the common ratio?

OpenStudy (timaashorty):

5 ?

OpenStudy (jdoe0001):

well.. \(\large \begin{array}{llll} 625&125&25&...\\ \hline\\ &\frac{625}{5}&\frac{125}{5} \end{array}\)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} a_1\\ \downarrow \\ 625&125&25&...\\ \hline\\ &625 \cdot \frac{1}{5}&125\cdot \frac{1}{5}\\ &&\qquad \ \uparrow \\ &&r=\textit{common ratio} \end{array}\qquad \bf \large a_n=a_1\cdot r^{n-1}\\ \quad \\ \Large a_{{\color{red}{ 10}}}=a_1\cdot r^{{\color{red}{ 10}}-1}\)

OpenStudy (timaashorty):

okay im kinda confused

OpenStudy (jdoe0001):

ok.... where?

OpenStudy (timaashorty):

do I plug in 1/5 in 'r'?

OpenStudy (jdoe0001):

yes, that's the common ratio in the geometric sequence notice that to get the next term, you just multiply the previous one by 1/5

OpenStudy (mathmale):

Notice that each successive term of the sequence 625,125,25,.. is 1/5 of the preceding term, so, yes, 1/5 is the common ratio. Can you now write a formula a sub n for the nth term of this sequence? Once you have that, you can easily predict the 10th term by letting n=10.

OpenStudy (timaashorty):

I got \[\frac{ 1 }{ 3125 }\] ?

OpenStudy (timaashorty):

is it A?

OpenStudy (timaashorty):

@jdoe0001 @mathmale

OpenStudy (mathmale):

Tima-Shorty, I did suggest this earlier: Notice that each successive term of the sequence 625,125,25,.. is 1/5 of the preceding term, so, yes, 1/5 is the common ratio. Can you now write a formula a sub n for the nth term of this sequence? Once you have that, you can easily predict the 10th term by letting n=10. Coming up with such a formula is an important skill in sequences and series.

OpenStudy (jdoe0001):

is it A? \(\large \checkmark\)

OpenStudy (mathmale):

With all due respect to y our wishes, Tima-Shorty, I'd much rather help you understand the underlying concept here, which is prediction of the nth term of a sequence. Twice you have not responded to my suggestion that you try writing a formula for that.